OpenStudy (anonymous):

The width w of a rectangular swimming pool is 7x2. The area A of the pool is 7x3 – 42x2. What is an expression for the length of the pool? Show your work.

OpenStudy (error1603):

(7x²) ( x - 6 ) ---------------------- 7x² L = x - 6

OpenStudy (error1603):

How are you today?

OpenStudy (anonymous):

i am good you

OpenStudy (error1603):

I'm great

OpenStudy (anonymous):

that;s good

OpenStudy (anonymous):

is there more i have to do?

OpenStudy (error1603):

No not really

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

can you help with another one?

OpenStudy (error1603):

This is the easy and fast way

OpenStudy (error1603):

sure

OpenStudy (anonymous):

As cars passed a checkpoint, the following speeds were clocked and recorded. Speed (mph): 55 62 61 54 68 72 59 61 70 16. What are the minimum, first quartile, median, third quartile, and maximum of the data set? Show your work.

OpenStudy (error1603):

First put the numbers in order: 54 55 59 61 61 62 68 70 72 The lowest number is the minimum. The highest is the maximum. Find the number that is in the middle (you can lightly cross out the outer numbers) The middle number is 61. That is the median. Then find the middle of each half. Since these numbers are between 2 numbers, you average the numbers (add then dividde by 2). 55+59=114/2=57 This is the first quartile. 68+70=138/2=69 This is the third quartile.

OpenStudy (anonymous):

Okay thank you so much

OpenStudy (error1603):

No problem